# Indicator Functions Detect Tangentially Transient Behaviour on Decaying Normally Hyperbolic Invariant Manifolds

### v1

@@ 4 Conclusions and Remarks
The decay of the NHIM and the related tangential transient effects can be seen in the restricted Poincaré map and also in the time delay function used as phase space structure indicator function. Now the reader may ask what are the advantages and disadvantages of these two possibilities to display the scenario as a function of the perturbation strength.
@@
The time delay function is easy and rapid to calculate and as any indicator function, it can be calculated and presented on any 2-dimensional surface in the constant energy manifold of the phase space. Moreover, it is not necessary to have detailed previous information on the location of the NHIM. The initial conditions leading to a high value of the time delay lie close to the stable manifold of localized subsets and guide us automatically towards the unstable invariant set.
@@
In contrast, the construction of the restricted Poincaré map needs the approximate knowledge of the position of the NHIM to start the calculations. As explained in detail in [17] so far we use a numerical method which constructs the restricted map and searches for the position of the NHIM in a combined form. As long as we have small perturbations and complete NHIMs without decay this method works quite well and is reliable. This method can display the finest details of the internal NHIM dynamics in graphical form including tangentially chaotic parts. However, for large perturbations and NHIMs in the process of decay, the method becomes more difficult. In some examples ( not yet published so far ) it has become extremely difficult to find transient parts of the NHIM in the process of decay by this method. It needed a lot of numerical effort and needed the special adaption of the method for each individual transient substructure of the NHIM.
@@
In this sense, the indicator function is the easier and more versatile method in general and it is also able to indicate the effects coming from the decaying parts of the NHIM. By relying on time delay functions exclusively one only loses some very fine details of the decay scenario. However, with the indicator function we only obtain information of the invariant sets that intersect the initial conditions chosen.
@@
We used an indicator function to display information on the localized chaotic invariant set of the system. This reminds us of the inverse chaotic scattering problem, where we use scattering functions and in particular also the time delay function as an indicator function to obtain information on the localized chaotic invariant set by asymptotic measurements. The only difference is that in scattering problems we use as the domain of the indicator function a subset of asymptotic initial conditions. In this sense, we move this domain to an infinite distance and for compensation, we have to subtract from the time delay the trivial asymptotic contributions which diverge to infinity. Otherwise, we have a perfect analogy. For general information on chaotic scattering see chapter 6 in [19], [30] , and chapter 4 in [31], for examples of the use of the time delay function in the inverse chaotic scattering problem see [32] and [33]. Our important new addition to this general basic idea archived in the present article is the special attention paid to tangential transient effects on unstable invariant subsets during their process of decay.

### v2

@@ 5 Conclusions and Remarks
The decay of the NHIM and the related tangential transient effects can be seen in the restricted Poincaré map and also in the time delay function used as a phase space structure indicator function. Now the reader may ask what are the advantages and disadvantages of these two possibilities to display the scenario as a function of the perturbation strength.
@@
The time delay function is easy and rapid to calculate and, it can be calculated and presented on any 2-dimensional surface in the constant energy manifold of the phase space. Moreover, it is not necessary to have detailed previous information on the location of the NHIM. The initial conditions leading to a high value of the time delay lie close to the stable manifold of localized subsets and guide us automatically towards the unstable invariant set.
@@
In contrast, the construction of the restricted Poincaré map needs the approximate knowledge of the position of the NHIM to start the calculations. As explained in detail in [18] so far we use a numerical method which constructs the restricted map and searches for the position of the NHIM $M_{\epsilon}$ in a combined form. As long as we have small perturbations and complete NHIMs without decay this method works quite well and is reliable. This method can display the finest details of the internal NHIM dynamics in graphical form including tangentially chaotic parts. However, for large perturbations and NHIMs in the process of decay, the method becomes more difficult. In some examples ( not yet published so far ) it has become extremely difficult to find transient parts of the NHIM in the process of decay by this method. It needed considerable numerical effort and a special adaptation of the method for each transient substructure of the NHIM.
@@
In this sense, the indicator function is the easier and more versatile method in general and it is also able to indicate the effects coming from the decaying parts of the NHIM. By relying on time delay functions exclusively one only loses some very fine details of the decay scenario. However, with the indicator function, we only obtain information on the invariant sets that intersect the initial conditions chosen.
@@
We used an indicator function to display information on the localized chaotic invariant set of the system. This reminds us of the inverse chaotic scattering problem, where we use scattering functions and in particular also the time delay function as an indicator function to obtain information on the localized chaotic invariant set by asymptotic measurements. The only difference is that in scattering problems we use as the domain of the indicator function a subset of asymptotic initial conditions. In this sense, we move this domain to an infinite distance and for compensation, we have to subtract from the time delay the trivial asymptotic contributions which diverge to infinity. Otherwise, we have a perfect analogy. For general information on chaotic scattering see chapter 6 in [20], [44] , and chapter 4 in [45]; for examples of the use of the time delay function in the inverse chaotic scattering problem see [46] and [47]. Our important new addition to this general basic idea achieved in the present article is the special attention paid to tangential transient effects on unstable invariant subsets during their process of decay.